# Computable Thermodynamic Formalism

### v1

@@ 2. Notation
Let $(X,d)$ be a metric space. We denote by $\mathcal{B}(X)$ the $\sigma$-algebra of all Borel subsets of $X$. For each subset $Y\subseteq X$, we denote the diameter of $Y$ by $\operatorname{diam}_{d}Y\coloneqq\sup\{d(x,y):x,\,y\in Y\}$, the interior of $Y$ by $\operatorname{int}Y$, and the characteristic function of $Y$ by $\mathbbm{1}_{Y}$.
@@
For a compact metric space $(X,d)$, we denote by $C(X)$ the space of continuous functions from $X$ to , and by $\mathcal{M}(X)$ (resp. $\mathcal{P}(X)$) the set of finite signed Borel measures (resp. Borel probability measures) on $X$. Let $g\colon X\rightarrow X$ be a Borel-measurable transformation. We denote by $\mathcal{M}(X,g)$ the set of $g$-invariant Borel probability measures on $X$. Moreover, for each Borel subset $C\in\mathcal{B}(X)$, $\mathcal{P}(X;C)$ denotes the set $\{\mu\in\mathcal{P}(X):\mu(C)=1\}$. By the Riesz representation theorem we can identify the dual of $C(X)$ with the space $\mathcal{M}(X)$. For $\mu\in\mathcal{M}(X)$, we use $\mathopen{\|}\mu\mathclose{\|}$ to denote the total variation norm of $\mu$, $\operatorname{supp}{\mu}$ to denote the support of $\mu$, and
@@ 3.1. Computable analysis
Computability over the reals. We begin by reviewing basic notations and concepts from classical recursion theory; for an introduction, see e.g. [Bri94, Section 3].
@@
Note that $\mathbb{N}^{k}$, for $k\in\mathbb{N}$, and $\mathbb{N}^{*}$ are all recursively enumerable sets by Definition 3.1. We then define partial recursive functions and recursive functions.
@@
Let $\{x_{i}\}_{i\in I}$ be a sequence of real numbers. We say that $\{x_{i}\}_{i\in I}$ is uniformly lower (resp. upper) semi-computable if there exist three recursive functions $f\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, $g\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, and $h\colon\mathbb{N}\times I\rightarrow\mathbb{N}$ such that for each $i\in I$, $\bigl\{(-1)^{h(n,i)}f(n,i)/g(n,i)\bigr\}_{n\in\mathbb{N}}$ is non-decreasing (resp. non-increasing) and converges to $x_{i}$ as $n\to+\infty$. Moreover, a real number $x$ is called lower (resp. upper) semi-computable if the sequence $\{x_{i}\}_{i\in\mathbb{N}}$ defined by $x_{i}\coloneqq x$ for each $i\in\mathbb{N}$ is uniformly lower (resp. upper) semi-computable.
@@
The above definition of a lower semi-computable open set differs slightly from the ones in [BBRY11, Definition 3.4] and [BRY14, Definition 2.4]. In our definition, we use extended ideal balls, which include the empty set.
The term recursively open set in the literature (e.g. [GHR11, Subsection 2.2 and Definition 2.4] and [HR09, Subsection 3.3]) is equivalent to the notion of lower semi-computable open set defined above. A detailed discussion of this equivalence is provided in [He25].
Note that we can algorithmically decide whether $s\in B$ for each ideal point $s\in\mathcal{S}$ and each extended ideal ball $B\in\mathcal{B}_{0}$. The following result then follows immediately from Definition 3.7.
@@
The following two results are two classical results in computable analysis which both follow immediately from Definitions 3.1 and 3.7 (see e.g. [He25]).
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$, and $x\in X$. We say that a function $\tau\colon\mathbb{N}\rightarrow\mathbb{N}$ is an oracle for $x$ if $\rho\bigl(s_{\tau(n)},x\bigr)<2^{-n}$ for each $n\in\mathbb{N}.$
@@
Let $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ be computable metric spaces with $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\{s^{\prime}_{n}\}_{n\in\mathbb{N}}$. Assume that $\{i_{n}\}_{n\in\mathbb{N}}$ is an effective enumeration of $I$, and $C_{i}\subseteq X$ for each $i\in I$. Then a sequence $\{f_{i}\}_{i\in I}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is called a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in I}$ if there exists an algorithm that, for all $m,\,n\in\mathbb{N}$, $x\in C_{i_{n}}$, and oracle $\tau$ for $x$, on input l,nl,\,n, and $\tau$, outputs $m\in\mathbb{N}$ with $\rho^{\prime}(s^{\prime}_{m},f_{i_{n}}(x))<2^{-l}$. We often omit the phrase “with respect to $\{C_{i}\}_{i\in I}$” when $C_{i}=X$ for all $i\in I$. Moreover, a function $f\colon X\rightarrow X^{\prime}$ is said to be a computable function on $C$ if $\{f_{i}\}_{i\in\mathbb{N}}$, defined by $f_{i}\coloneqq f$ for all $i\in\mathbb{N}$, is a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in\mathbb{N}}$ defined by $C_{i}\coloneqq C$ for all $i\in\mathbb{N}$. We often omit the phrase “with respect to CC” when $C=X$.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $C_{i}\subseteq X$ for each $i\in I$. A sequence $\{f_{i}\}_{i\in I}$ of functions $f_{i}\colon X\rightarrow\mathbb{R}$ is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in I}$ if there exists a sequence $\{g_{i,n}\}_{(i,n)\in I\times\mathbb{N}}$ of uniformly computable functions $g_{i,n}\colon X\rightarrow\mathbb{R}$ with respect to $\{C_{i}\}_{i\in I}$ such that for each $i\in I$ and each $x\in C_{i}$, $\{g_{i,n}(x)\}_{n\in\mathbb{N}}$ is non-increasing (resp. non-decreasing) and converges to $f_{i}(x)$ as $n\to+\infty$. We often omit the phrase “with respect to $\{C_{i}\}_{i\in I}$” when $C_{i}=X$ for each $i\in I$. Moreover, a function $f\colon X\rightarrow\mathbb{R}$ is said to be an upper (resp. a lower) semi-computable function on $C$ if $\{f_{i}\}_{i\in\mathbb{N}}$ defined by $f_{i}\coloneqq f$ for each $i\in\mathbb{N}$, is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in\mathbb{N}}$ defined by $C_{i}\coloneqq C$ for all $i\in\mathbb{N}$. We often omit the phrase “with respect to CC” when $C=X$.
The following proposition is an immediate consequence of Proposition 3.14 (see e.g. [He25]).
@@
The following result indicates that the characteristic function of a lower semi-computable open set is a lower semi-computable function (see e.g. [He25]).
@@
    For each $x\in X$, $\{h_{n,i}(x)\}_{n\in\mathbb{N}}$ is non-decreasing and converges to $\mathbbm{1}_{U_{i}}(x)$ as $n\to+\infty$.
@@
We collect some fundamental properties of recursively compact sets (cf. [GHR11, Propositions 1 & 3]; see also [He25]).
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Suppose $X$ is recursively compact, $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, and $\{K_{i,h}\}_{(i,h)\in L}$ is uniformly recursively compact. Define $L_{h}\coloneqq\{(i,h)\in L:i\in I\}$ for each $h\in H$. Then the following statements are true:
@@
    If the function $F\colon H\rightarrow\mathbb{N}$ defined by $F(h)\coloneqq\operatorname{card}\mathopen{(}L_{h}\mathclose{)}$ for $h\in H$ is recursive, then $\{\bigcup\{K_{i,h}:(i,h)\in L_{h}\}\}_{h\in H}$ is uniformly recursively compact.
@@
Moreover, given the recursive compactness of $X$, the computability of functions is preserved under a finite number of operations among additions and multiplications. We summarize this property in the following result (cf. [We00, Corollary 4.3.4]; see also [He25]).
@@
Finally, we record the following useful characterization of complete recursively precompact metric spaces (cf. [GHR11, Proposition 4]; see also [He25]).
@@
Computability of probability measures. Building upon the theory of computable functions and recursively compact sets, we now discuss the computability of probability measures. We begin by reviewing the computable structure on the measure space $\mathcal{P}(X)$ introduced in [HR09, Section 4] (cf. [HR09, Proposition 4.1.3]; see also [He25]).
@@
We now recall a key result on the computability of the integration function (cf. [HR09, Corollary 4.3.2]; see also [He25]).
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact. Assume that $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, $\{U_{i,h}\}_{(i,h)\in L}$ is uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$, and $\{r_{i,h}\}_{(i,h)\in L}$ is a sequence of uniformly computable real numbers. Then $\{\mathcal{K}_{i}\}_{i\in I}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$. Here $\mathcal{K}_{i}\coloneqq\{\mu\in\mathcal{P}(X):\mu(U_{i,h})\leqslant r_{i,h}\text{ for each }(i,h)\in L_{i}\}$, and $L_{i}\coloneqq\{(i,h)\in L:h\in H\}$ for each $i\in I$.
Recall the definition of $\mathcal{M}(X,T;Y)$ from (1.1). The following result indicates the recursive compactness of the set $\mathcal{M}(X,T;Y)$ (cf. [BHLZ25, Lemma 4.12]; see also [He25]).
@@
The following result is useful in the proof of the main result of this article (see e.g. [He25]).
@@ 3.2. Thermodynamic formalism
Assume that $T\colon X\rightarrow X$ is a measure-preserving transformation of $(X,\mathcal{B},\mu)$. Consider a partition $\xi=\{A_{h}:h\in H\}$ of $X$. For each $n\in\mathbb{N}$, $T^{-n}(\xi)$ denotes the partition $\bigl\{T^{-1}(A_{h}):h\in H\bigr\}$, and $\xi^{n}_{T}$ denotes the join $\xi\vee T^{-1}(\xi)\vee\cdots\vee T^{-(n-1)}(\xi)$. The entropy of $\xi$ is $H_{\mu}(\xi)\coloneqq-\sum_{h\in H}\mu(A_{h})\log(\mu(A_{h}))\in[0,+\infty]$, where $0\log 0$ is defined to be zero. One can show (see e.g. [Wa82, Chapter 4]) that if $H_{\mu}(\xi)<+\infty$, then $\lim\limits_{n\to+\infty}H_{\mu}(\xi^{n}_{T})/n$ exists. We denote this limit by $h_{\mu}(T,\xi)$ and call it the measure-theoretic entropy of $T$ relative to $\xi$. The measure-theoretic entropy of $T$ for $\mu$ is defined as
@@
4. Approach I
In this section, we establish the computability of equilibrium states for certain dynamical systems with upper semi-continuous measure-theoretic entropy functions. We begin by recalling several definitions and results from functional analysis: a characterization of the set ${C(X)}^{*}_{\phi,P_{T}}$ in (4.1) and its relation to the set $\mathcal{E}(T,\phi)$ of equilibrium states in (4.2). We then apply these results to verify the recursive compactness of ${C(X)}^{*}_{\phi,P_{T}}$, thereby completing the proof of Theorem 1.1.
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ a continuous map with finite topological entropy. The measure-theoretic entropy function of $T$ is the function $\mu\mapsto h_{\mu}(T)$ (defined in (3.1)) on the space $\mathcal{M}(X,T)$ of $T$-invariant Borel probability measures, where $\mathcal{M}(X,T)$ is equipped with the weak topology. The topological pressure function of $T$, denoted by $P_{T}$, is the function $\phi\mapsto P(T,\phi)$ (defined in (3.2)) on $C(X)$. For each $\mu\in\mathcal{M}(X,T)$, we denote by $\overline{h}_{\mu}(T)$ the upper semi-continuous regularization of the measure-theoretic entropy function. More precisely, $\overline{h}_{\mu}(T)$ is the supremum of all limit suprema $\limsup_{n\to\infty}h_{\mu_{n}}(T)$, where $\{\mu_{n}\}_{n\in\mathbb{N}}$ ranges over all sequences in $\mathcal{M}(X,T)$ that converge to $\mu$ in the weak topology.

### v2

@@ 2. Notation
Let $(X,d)$ be a metric space. We denote by $\mathcal{B}(X)$ the $\sigma$-algebra of all Borel subsets of $X$. For each subset $Y\subseteq X$, we denote the diameter of $Y$ by $\operatorname{diam}_{d}Y\coloneqq\sup\{d(x,y):x,\,y\in Y\}$, the interior of $Y$ by $\operatorname{int}_{\circ}Y$, and the characteristic function of $Y$ by $\mathbbm{1}_{Y}$.
@@
For a compact metric space $(X,d)$, we denote by $C(X)$ the space of continuous functions from $X$ to , and by $\mathcal{M}(X)$ (resp. $\mathcal{P}(X)$) the set of finite signed Borel measures (resp. Borel probability measures) on $X$. Let $g\colon X\rightarrow X$ be a Borel-measurable transformation. We denote by $\mathcal{M}(X,g)$ the set of $g$-invariant Borel probability measures on $X$. Moreover, for each Borel subset $C\in\mathcal{B}(X)$, $\mathcal{P}(X;C)$ denotes the set $\{\mu\in\mathcal{P}(X):\mu(C)=1\}$. By the Riesz representation theorem, we can identify the dual of $C(X)$ with the space $\mathcal{M}(X)$. For $\mu\in\mathcal{M}(X)$, we use $\mathopen{\|}\mu\mathclose{\|}$ to denote the total variation norm of $\mu$, $\operatorname{supp}{\mu}$ to denote the support of $\mu$, and
@@ 3.1. Computable analysis
Computability over the reals. We begin by reviewing basic notations and concepts from classical recursion theory; for an introduction, see e.g. [Bri94, Chapter 3].
@@
Note that $\mathbb{N}^{k}$, for $k\in\mathbb{N}$, and $\mathbb{N}^{*}$ are all recursively enumerable sets by Definition 3.1.
@@
Let $\{x_{i}\}_{i\in I}$ be a sequence of real numbers. We say that $\{x_{i}\}_{i\in I}$ is uniformly lower (resp. upper) semi-computable if there exist three recursive functions $f\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, $g\colon\mathbb{N}\times I\rightarrow\mathbb{N}$, and $h\colon\mathbb{N}\times I\rightarrow\mathbb{N}$ such that for each $i\in I$, $\bigl\{(-1)^{h(n,i)}f(n,i)/g(n,i)\bigr\}_{n\in\mathbb{N}}$ is nondecreasing (resp. nonincreasing) and converges to $x_{i}$ as $n\to+\infty$. Moreover, a real number $x$ is called lower (resp. upper) semi-computable if the sequence $\{x_{i}\}_{i\in\mathbb{N}}$ defined by $x_{i}\coloneqq x$ for each $i\in\mathbb{N}$ is uniformly lower (resp. upper) semi-computable.
@@
The above definition of a lower semi-computable open set differs slightly from the ones in [BBRY11, Definition 3.4] and [BRY14, Definition 2.4]. In our definition, we use extended ideal balls, which include the empty set $\emptyset$.
The term recursively open set in the literature (e.g. [GHR11, Subsection 2.2 and Definition 2.4] and [HR09, Subsection 3.3]) is equivalent to the notion of lower semi-computable open set defined above. A detailed discussion of this equivalence is provided in [He25, Subsection 3.3].
Note that we can algorithmically decide whether $s\in B$ for each ideal point $s\in\mathcal{S}$ and each extended ideal ball $B\in\mathcal{B}_{0}$. The following result then follows immediately from Definition 3.7 (see e.g. [He25, Proposition 3.9]).
@@
The following two results are two classical results in computable analysis which both follow immediately from Definitions 3.1 and 3.7 (see e.g. [He25, Propositions 3.10 & 3.11]).
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space with $\mathcal{S}=\{s_{i}\}_{i\in\mathbb{N}}$, and $x\in X$. We say that a function $\tau\colon\mathbb{N}\rightarrow\mathbb{N}$ is an oracle for $x$ if $\rho(s_{\tau(n)},x)<2^{-n}$ for each $n\in\mathbb{N}.$
@@
Let $(X,\,\rho,\,\mathcal{S})$ and $(X^{\prime},\,\rho^{\prime},\,\mathcal{S}^{\prime})$ be computable metric spaces with $\mathcal{S}=\{s_{n}\}_{n\in\mathbb{N}}$ and $\mathcal{S}^{\prime}=\{s^{\prime}_{n}\}_{n\in\mathbb{N}}$. Assume that $\{i_{n}\}_{n\in\mathbb{N}}$ is an effective enumeration of $I$, and $C_{i}\subseteq X$ for each $i\in I$. Then a sequence $\{f_{i}\}_{i\in I}$ of functions $f_{i}\colon X\rightarrow X^{\prime}$ is called a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in I}$ if there exists an algorithm that, for all $l,\,n\in\mathbb{N}$, $x\in C_{i_{n}}$, and oracle $\tau$ for $x$, on input l,nl,\,n, and $\tau$, outputs $m\in\mathbb{N}$ with $\rho^{\prime}(s^{\prime}_{m},f_{i_{n}}(x))<2^{-l}$. We often omit the phrase “with respect to $\{C_{i}\}_{i\in I}$” when $C_{i}=X$ for all $i\in I$. Moreover, a function $f\colon X\rightarrow X^{\prime}$ is said to be a computable function on $C$ if $\{f_{i}\}_{i\in\mathbb{N}}$, defined by $f_{i}\coloneqq f$ for all $i\in\mathbb{N}$, is a sequence of uniformly computable functions with respect to $\{C_{i}\}_{i\in\mathbb{N}}$ defined by $C_{i}\coloneqq C$ for all $i\in\mathbb{N}$. We often omit the phrase “with respect to CC” when $C=X$.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, $\{i_{n}\}_{n\in\mathbb{N}}$ be an effective enumeration of $I$, and $C_{i}\subseteq X$ for each $i\in I$. A sequence $\{f_{i}\}_{i\in I}$ of functions $f_{i}\colon X\rightarrow\mathbb{R}$ is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in I}$ if there exists an algorithm that, for all $l,\,n\in\mathbb{N}$, $x\in C_{i_{n}}$, and oracle $\tau$ for $x$, on input l,nl,\,n, and $\tau$, outputs $q_{l,n,\tau}\in\mathbb{Q}$ such that for each $n\in\mathbb{N}$, each $x\in C_{i_{n}}$, and each oracle $\tau$ for $x$, $\{q_{l,n,\tau}\}_{l\in\mathbb{N}}$ is nonincreasing (resp. nondecreasing) and converges to $f_{i_{n}}(x)$ as $l\to+\infty$. We often omit the phrase “with respect to $\{C_{i}\}_{i\in I}$” when $C_{i}=X$ for each $i\in I$. Moreover, a function $f\colon X\rightarrow\mathbb{R}$ is said to be an upper (resp. a lower) semi-computable function on $C$ if $\{f_{i}\}_{i\in\mathbb{N}}$ defined by $f_{i}\coloneqq f$ for each $i\in\mathbb{N}$, is a sequence of uniformly upper (resp. lower) semi-computable functions with respect to $\{C_{i}\}_{i\in\mathbb{N}}$ defined by $C_{i}\coloneqq C$ for all $i\in\mathbb{N}$. We often omit the phrase “with respect to CC” when $C=X$.
The following proposition is an immediate consequence of Proposition 3.14 (see e.g. [He25, Theorem 3.19]).
@@
The following result indicates that the characteristic function of a lower semi-computable open set is a lower semi-computable function (see e.g. [He25, Proposition 3.33]).
@@
    For each $x\in X$, $\{h_{n,i}(x)\}_{n\in\mathbb{N}}$ is nondecreasing and converges to $\mathbbm{1}_{U_{i}}(x)$ as $n\to+\infty$.
@@
We collect some fundamental properties of recursively compact sets (cf. [GHR11, Propositions 1 & 3]; see also [He25, Proposition 3.23]).
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Suppose $X$ is recursively compact, $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, and $\{K_{i,h}\}_{(i,h)\in L}$ is uniformly recursively compact. Denote $L_{h}\coloneqq\{(i,h)\in L:i\in I\}$ for each $h\in H$. Then the following statements are true:
@@
    If the function $F\colon H\rightarrow\mathbb{N}$ defined by $F(h)\coloneqq\operatorname{card}L_{h}$ for $h\in H$ is recursive, then $\{\bigcup\{K_{i,h}:(i,h)\in L_{h}\}\}_{h\in H}$ is uniformly recursively compact.
@@
Moreover, given the recursive compactness of $X$, the computability of functions is preserved under a finite number of operations among additions and multiplications. We summarize this property in the following result (cf. [We00, Corollary 4.3.4]; see also [He25, Proposition 3.26]).
@@
Finally, we record the following useful characterization of complete recursively precompact metric spaces (see e.g. [GHR11, Proposition 4]).
@@
Computability of probability measures. Building upon the theory of computable functions and recursively compact sets, we now discuss the computability of probability measures. We begin by reviewing the computable structure on the measure space $\mathcal{P}(X)$ introduced in [HR09, Section 4] (cf. [HR09, Proposition 4.1.3]; see also [He25, Proposition 3.29]).
@@
We now recall a key result on the computability of the integration function (cf. [HR09, Corollary 4.3.2]; see also [He25, Proposition 3.30]).
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact. Assume that $H$ and $L$ are two nonempty recursively enumerable sets with $L\subseteq I\times H$, $\{U_{i,h}\}_{(i,h)\in L}$ is uniformly lower semi-computable open in $(X,\,\rho,\,\mathcal{S})$, and $\{r_{i,h}\}_{(i,h)\in L}$ is a sequence of uniformly computable real numbers. Define, for each $i\in I$, $L_{i}\coloneqq\{(i,h)\in L:h\in H\}$ and $\mathcal{K}_{i}\coloneqq\{\mu\in\mathcal{P}(X):\mu(U_{i,h})\leqslant r_{i,h}\text{ for each }(i,h)\in L_{i}\}$. Then $\{\mathcal{K}_{i}\}_{i\in I}$ is uniformly recursively compact in $(\mathcal{P}(X),\,W_{\rho},\,\mathcal{Q}_{\mathcal{S}})$.
Recall the definition of $\mathcal{M}(X,T;Y)$ from (1.1). The following result indicates the recursive compactness of the set $\mathcal{M}(X,T;Y)$ (cf. [BHLZ25, Lemma 4.12]; see also [He25, Theorem 3.36]).
@@
The following result is useful in the proof of the main result of this article (see e.g. [He25, Proposition 3.35]).
@@ 3.2. Thermodynamic formalism
Assume that $T\colon X\rightarrow X$ is a measure-preserving transformation of $(X,\mathcal{B},\mu)$. Consider a partition $\xi=\{A_{h}:h\in H\}$ of $X$. For each $n\in\mathbb{N}$, $T^{-n}(\xi)$ denotes the partition $\bigl\{T^{-1}(A_{h}):h\in H\bigr\}$, and $\xi^{n}_{T}$ denotes the join $\xi\vee T^{-1}(\xi)\vee\cdots\vee T^{-(n-1)}(\xi)$. The entropy of $\xi$ is $H_{\mu}(\xi)\coloneqq-\sum_{h\in H}\mu(A_{h})\log(\mu(A_{h}))\in[0,+\infty]$, where $0\log 0$ is defined to be zero. One can show that if $H_{\mu}(\xi)<+\infty$, then $\lim\limits_{n\to+\infty}H_{\mu}(\xi^{n}_{T})/n$ exists (see e.g. [Wa82, Chapter 4]). We denote this limit by $h_{\mu}(T,\xi)$ and call it the measure-theoretic entropy of $T$ relative to $\xi$. The measure-theoretic entropy of $T$ for $\mu$ is defined as
@@
4. Approach I: Global approach via convex analysis
In this section, we establish the computability of equilibrium states for certain dynamical systems with upper semicontinuous measure-theoretic entropy functions. We begin by recalling several definitions and results from functional analysis: a characterization of the set ${C(X)}^{*}_{\phi,P_{T}}$ in (4.1) and its relation to the set $\mathcal{E}(T,\phi)$ of equilibrium states in (4.2). We then apply these results to verify the recursive compactness of ${C(X)}^{*}_{\phi,P_{T}}$, thereby completing the proof of Theorem 1.1.
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ a continuous map with finite topological entropy. The measure-theoretic entropy function of $T$ is the function $\mu\mapsto h_{\mu}(T)$ (defined in (3.1)) on the space $\mathcal{M}(X,T)$ of $T$-invariant Borel probability measures, where $\mathcal{M}(X,T)$ is equipped with the weak topology. The topological pressure function of $T$, denoted by $P_{T}$, is the function $\phi\mapsto P(T,\phi)$ (defined in (3.2)) on $C(X)$. For each $\mu\in\mathcal{M}(X,T)$, we denote by $\overline{h}_{\mu}(T)$ the upper semicontinuous regularization of the measure-theoretic entropy function. More precisely, $\overline{h}_{\mu}(T)$ is the supremum of all limit suprema $\limsup_{n\to\infty}h_{\mu_{n}}(T)$, where $\{\mu_{n}\}_{n\in\mathbb{N}}$ ranges over all sequences in $\mathcal{M}(X,T)$ that converge to $\mu$ in the weak topology.